### v2

# UNITAL OPERATOR SPACES AND DISCRETE GROUPS

Abstract.
The first result regards reflexive unital masa bimodules acting on a Hilbert space. We show that any such space can be written as an intersection of terms $\mathbb{A}+\mathbb{B}$, $\mathbb{A}$ and $\mathbb{B}$ being CSL algebras. Next, we introduce the trivial intersection property for concrete operator spaces and we show that a unital space with this property has no nontrivial boundary ideals. We provide various examples of spaces with this property, such as strongly reflexive masa bimodules and completely distributive CSL algebras. We show that unital operator spaces acting on $\ell^{2}(\Gamma)$ for any set $\Gamma$, that contain the masa $\ell^{\infty}(\Gamma)$, possess the trivial intersection property, and we use this to prove that a unital surjective complete isometry between such spaces is actually a unitary equivalence. Then, we apply these results to $w^{*}$-closed $\ell^{\infty}(G)$-bimodules acting on $\ell^{2}(G)$ for a group $G$ and we relate them to algebraic properties of $G$.
Key words and phrases:
masa bimodules, unital operator spaces, operator systems, reflexive spaces, CSL algebras, $C^{*}$-envelopes
1991 Mathematics Subject Classification:
Primary: 46L07, 47L35; Secondary: 47A15, 47L25
This work was financially supported by the ”Andreas Mentzelopoulos Foundation”.
1. Introduction and Preliminaries
In this note, our spaces of interest are concrete unital operator spaces, i.e. norm-closed subspaces $\mathcal{U}\subseteq B(\mathcal{H})$ that contain the unit of $B(\mathcal{H})$. Here, $B(\mathcal{H})$ denotes the space of all bounded linear operators acting on a Hilbert space $\mathcal{H}$.
In Section 2, we deal with reflexive unital masa bimodules, i.e. reflexive unital operator spaces $\mathcal{U}$ that satisfy $\mathbb{D}\mathcal{U}\mathbb{D}\subseteq\mathcal{U}$ for some masa (maximal abelian self-adjoint algebra) $\mathbb{D}\subseteq B(\mathcal{H})$. The main result is Corollary 2.3, where we prove an explicit expression for any such space as an intersection of reflexive spaces that can be written as sums of two CSL algebras. In case $\mathcal{U}$ is also self-adjoint (i.e. an operator system), the forementioned expression can be modified so that it reflects the fact $\mathcal{U}^{*}=\mathcal{U}$.
In Section 3, we focus our attention on concrete unital operator spaces and we introduce a property called the trivial intersection property. It follows directly that every unital subspace $\mathcal{U}\subseteq B(\mathcal{H})$ with this property has no boundary ideals, hence its $C^{*}$-envelope $C^{*}_{e}(\mathcal{U})$ coincides with $C^{*}(\mathcal{U})$, the $C^{*}$-algebra $\mathcal{U}$ generates in $B(\mathcal{H})$. It turns out that this property appears in several different cases. To name a few, every strongly reflexive masa bimodule has the trivial intersection property, as well as every completely distributive CSL algebra. We also show that every subspace of $B(\ell^{2}(\Gamma))$ that contains the masa $\ell^{\infty}(\Gamma)$ has the trivial intersection property and we use this to prove Theorem 3.10, which states that every unital surjective complete isometry between unital spaces that contain a discrete masa is a unitary equivalence.
Finally, in Section 4, we apply results from Section 3 to subspaces of $B(\ell^{2}(G))$, where $G$ is a group. We are interested in spaces of the form $\mathbb{M}(\Omega)$ (see (5)), i.e. the space of all operators supported in a set $\Omega\subseteq G\times G$. The support of an operator was first introduced by Arveson in [Arv74] and generalized by Erdos, Katavolos and Shulman in [EKS98], and our definition for $\mathbb{M}(\Omega)$ is based on their work. In fact, we mostly deal with sets of the form $E^{\star}$ for $E\subseteq G$, where $E^{\star}=\{(g,h)\in G\times G:g^{-1}h\in E\}$ and we investigate how algebraic properties of $E$ are related to algebraic and topological properties of $\mathbb{M}(E^{\star})$.
Arveson introduced in [Arv74] the notion of a Commutative Subspace Lattice (CSL). By definition, a CSL is a complete sublattice of projections acting on a Hilbert space that commute with each other and which contains the zero and the identity operator. A CSL algebra is any algebra of the form $(\mathcal{L})$ [^1] For any set $\mathcal{L}$ of projections, Alg():=$(\mathcal{L}):=\{T\in B(\mathcal{H}):\ TP\subseteq P\ \ \text{for each}\ \ P\in\mathcal{L}\}$., where $\mathcal{L}$ is a CSL. It is well-known that $\mathscr{A}\subseteq B(\mathcal{H})$ is a CSL algebra if and only if it is a reflexive algebra (in the sense of Halmos [Hal71]) that contains a masa. As a generalization of CSL algebras, masa bimodules have been widely studied by Arveson, Davidson, Eleftherakis, Erdos, Katavolos, Ludwig, Shulman, Todorov, Turowska et al. (see e.g. [Arv74], [Dav88], [EKS98], [Ele14], [LT06], [ST04], [STT20], [Tod02], [TT16]).
The definition of the $C^{*}$-envelope, as well as its universal property is thanks to M. Hamana [Ham79], although W. Arveson had already proved their existence and the universal property in several cases [Arv69, Arv72]. For an explicit definition of the $C^{*}$-envelope and description of the universal property, the reader is referred to [BM04, Theorem 4.3.1]. We shall need the fact that for every concrete unital operator space $\mathcal{U}\subseteq B(\mathcal{H})$, $C^{*}_{e}(\mathcal{U})$ is (*-isomorphic to) $C^{*}(\mathcal{U})/J$ where $J$ (called the Shilov boundary of $\mathcal{U}$) is the largest closed two-sided ideal of $C^{*}(\mathcal{U})$ with the property that the canonical map
$q\circ i:\mathcal{U}\rightarrow C^{*}(\mathcal{U})/J$
is a complete isometry, where $i:\mathcal{U}\rightarrow C^{*}(\mathcal{U})$ is the inclusion map and $q$ is the quotient map from $C^{*}(\mathcal{U})$ onto $C^{*}(\mathcal{U})/J$ (any closed two sided ideal of $C^{*}(\mathcal{U})$ with this property is called a boundary ideal for $\mathcal{U}$). For a proof of the above fact, see e.g. [BM04, 4.3.2]. For a further discussion on (unital) operator spaces, see [BM04], [ER00], [Pau02] or [Pis03].
As a final clarification, any element $(a_{\gamma})_{\gamma\in\Gamma}\in\ell^{\infty}(\Gamma)$ induces an associated operator in $B(\ell^{2}(\Gamma))$, namely
$(x_{\gamma})_{\gamma\in\Gamma}\mapsto(a_{\gamma}x_{\gamma})_{\gamma\in\Gamma}.$
So, by $\ell^{\infty}(\Gamma)$ we mean all operators of this kind. It is not so difficult to see that this is actually a masa acting on $\ell^{2}(\Gamma)$.
In all Sections, all subspaces are assumed to be norm-closed.
2. An expression for reflexive unital masa bimodules in terms of CSL algebras
In this section, we use several concepts introduced by J. A. Erdos in [Erd86], so we shall briefly discuss what was introduced there. We use the letters $\mathcal{H}$, $\mathcal{H}_{1}$ etc. to denote a Hilbert space.
For Hilbert spaces $\mathcal{H}_{1}$, $\mathcal{H}_{2}$ the reflexive hull of a subset $\mathcal{U}\subseteq B(\mathcal{H}_{1},\mathcal{H}_{2})$ is defined as
$\text{Ref}(\mathcal{U}):=\{T\in B(\mathcal{H}_{1},\mathcal{H}_{2})\ :\ Tx\in\overline{[\mathcal{U}x]}\ \text{for each}\ x\in\mathcal{H}_{1}\}.$
Evidently, $\mathcal{U}\subseteq$ Ref$(\mathcal{U})$ and $\mathcal{U}$ is called reflexive if $\mathcal{U}=$ Ref$(\mathcal{U})$. This generalizes the definition of a Reflexive Algebra, introduced by Halmos in [Hal71]. Erdos proved [Erd86, Theorem 4.2 (i)] that the two definitions are equivalent in case $\mathcal{U}\subseteq B(\mathcal{H})$ is a unital algebra.
We denote the set of all projections contained in a set $V\subseteq B(\mathcal{H})$ by $\mathcal{P}(V)$ (we sometimes identify a projection with its range). Then, given any subset $\mathcal{U}\subseteq B(\mathcal{H}_{1},\mathcal{H}_{2})$, we obtain a map
$\phi=\text{Map}(\mathcal{U}):\mathcal{P}(B(\mathcal{H}_{1}))\rightarrow\mathcal{P}(B(\mathcal{H}_{2}))$
$P\mapsto\overline{[\mathcal{U}(P)]}.$
We have the following Lemma:
Lemma 2.1.
Let $\mathcal{U}$ be a $\mathbb{D}$-bimodule for a masa $\mathbb{D}\subseteq B(\mathcal{H})$. Then, the image of $\phi=\textup{Map}(\mathcal{U})$ is contained in $\mathcal{P}(\mathbb{D})$. Moreover,
(1) $\textup{Ref}(\mathcal{U})=\{T\in B(\mathcal{H}):\phi(P)^{\perp}TP=0\ \text{foreach}\ P\in\mathbb{D}\}.$
If $\mathcal{H}$ is separable, there exists a sequence of projections $P_{n}$ in $\mathbb{D}$, such that
(2) $\textup{Ref}(\mathcal{U})=\{T\in B(\mathcal{H}):\phi(P_{n})^{\perp}TP_{n}=0\ \text{for each}\ n\}.$
Proof.
It is trivial to check that for every $T\in\mathbb{D}$ and $P\in\mathcal{P}(B(\mathcal{H}))$ we have $T\phi(P)\subseteq\phi(P)$ and $T\phi(P)^{\perp}\subseteq\phi(P)^{\perp}$, which implies that $T\phi(P)=\phi(P)T$, thus $\phi(P)\in\mathcal{P}(\mathbb{D})$.
By a straightforward verification, one observes that
$\text{Ref}(\mathcal{U})\subseteq\{T\in B(\mathcal{H}):\phi(P)^{\perp}TP=0\ \text{for each}\ P\in\mathbb{D}\}.$
For the converse, let $\xi\in\mathcal{H}$. It is easy to see that the projection $P=\overline{\mathbb{D}\xi}\in\mathcal{P}(\mathbb{D})$, thus
$T\xi\in TP\subseteq\phi(P)=\overline{\mathcal{U}\xi}$
hence $T\in$ Ref$(\mathcal{U})$ by definition.
If we assume in addition that $\mathcal{H}$ is separable, then the predual $B(\mathcal{H}\oplus\mathcal{H})_{*}$ is also separable, which implies that the unit ball $\mathcal{B}_{1}$ of $B(\mathcal{H}\oplus\mathcal{H})$ is $w^{*}$-metrizable. By the Banach-Alaoglu Theorem, $\mathcal{B}_{1}$ is also $w^{*}$-compact, thus separable. We define
:=$\mathcal{B}:=\{\phi(P)\oplus P:P\in\mathcal{P}(\mathbb{D})\}\subseteq\mathcal{B}_{1}.$
By our earlier observations, $\mathcal{B}$ is also separable (in the subspace $w^{*}$-topology), thus there exists a sequence $(P_{n})\subseteq\mathcal{P}(\mathbb{D})$, such that the set $\{\phi(P_{n})\oplus P_{n}:n\in\mathbb{N}\}$ is $w^{*}$-dense in $\mathcal{B}$. Now, (2) follows easily from (1). ∎
For the rest of this section, $\mathbb{D}$ denotes a masa acting on a Hilbert space $\mathcal{H}$. For a reflexive unital $\mathbb{D}$-bimodule $\mathcal{U}\subseteq B(\mathcal{H})$ and and for ϕ:=assignitalic-$\phi:=$ Map$(\mathcal{U})$, we clearly have $\mathbb{D}\subseteq\mathcal{U}$. We also note that since $\mathcal{U}$ is unital,
$\phi(P)\geq P\ \ \text{for each}\ \ P\in\mathcal{P}(B(\mathcal{H})).$
For a subspace $\mathcal{V}\subseteq B(\mathcal{H})$, we define the diagonal of $\mathcal{V}$ as $\Delta(\mathcal{V}):=\mathcal{V}\cap\mathcal{V}^{*}$.
Proposition 2.2.
Suppose that $\mathcal{U}\subseteq B(\mathcal{H})$ is a reflexive unital $\mathbb{D}$-bimodule. For $\phi=\textup{Map}(\mathcal{U})$, $P\in\mathbb{D}$ and
$X:=\{T\in B(\mathcal{H})\ :\ \phi(P)^{\perp}TP=0\}$
the following hold:
  • (i)
    $X=\mathbb{A}_{1}+\mathbb{A}_{2}$ for certain CSL algebras $\mathbb{A}_{1}$, $\mathbb{A}_{2}$.
  • (ii)
    If in addition $\mathcal{U}$ is self-adjoint, we can write
    $\Delta(X)=\mathbb{B}+\mathbb{B}^{*}$
    for a certain CSL algebra $\mathbb{B}$.
Proof.
We first prove $(i)$. We define the projection $Q:=\phi(P)-P$ and we set
$\mathbb{A}_{1}:=\{T\in X:\ Q^{\perp}TQ=0\}=\text{Alg}(\{Q\})\cap X$
$\mathbb{A}_{2}:=\{T\in X:\ QTQ^{\perp}=0\}=\text{Alg}(\{Q^{\perp}\})\cap X.$
Since $Q\in\mathbb{D}$, it is clear that $X=\mathbb{A}_{1}+\mathbb{A}_{2}$. In order to show that $\mathbb{A}_{1}$ and $\mathbb{A}_{2}$ are algebras, the only nontrivial part is to prove that $\mathbb{A}_{i}\mathbb{A}_{i}\subseteq X$ for $i=1,2$. One can easily verify that $XQ^{\perp}X\subseteq X$, and that
$\mathbb{A}_{1}Q\mathbb{A}_{1}=Q\mathbb{A}_{1}Q\mathbb{A}_{1}\ \ \text{and}\ \ \mathbb{A}_{2}Q\mathbb{A}_{2}=\mathbb{A}_{2}Q\mathbb{A}_{2}Q$
which implies $\mathbb{A}_{i}Q\mathbb{A}_{i}\subseteq X$ for $i=1,2$, since $Q\perp\phi(P)^{\perp}$ and $Q\perp P$. Since $\mathbb{A}_{1}$, $\mathbb{A}_{2}$ are clearly reflexive and they both contain $\mathbb{D}$, they are CSL algebras.
Now, we prove $(ii)$. Clearly,
$\Delta(X)=\{T\in B(\mathcal{H})\ :\ \phi(P)^{\perp}TP=PT\phi(P)^{\perp}=0\}.$
We define
$\mathbb{B}:=\{T\in B(\mathcal{H})\ :\ P^{\perp}TP=\phi(P)T\phi(P)^{\perp}=0\}=\text{Alg}(\{P,\phi(P)^{\perp}\}).$
Since $P\leq\phi(P)$, $\mathbb{B}\subseteq\Delta(X)$ and since $\Delta(X)$ is self-adjoint, we have $\mathbb{B}+\mathbb{B}^{*}\subseteq\Delta(X)$. For the converse, for $Q=\phi(P)-P$ (as in $(i)$), a straightforward verification shows that $Q\Delta(X)Q$, $Q^{\perp}\Delta(X)Q$ and $Q^{\perp}\Delta(X)Q^{\perp}$ are all contained in $\mathbb{B}$, therefore $Q\Delta(X)Q^{\perp}\subseteq\mathbb{B}^{*}$ and we have proved
$\Delta(X)=\mathbb{B}+\mathbb{B}^{*}.$
By the same arguments as in $(i)$, $\mathbb{B}$ is a CSL algebra.
Corollary 2.3.
Suppose that $\mathcal{H}$ is a separable Hilbert space and that $\mathcal{U}\subseteq B(\mathcal{H})$ is a reflexive unital $\mathbb{D}$-bimodule.
  • (i)
    We can write $\mathcal{U}$ in the form
    (3) $\mathcal{U}=\bigcap_{n=1}^{\infty}(\mathbb{A}_{1,n}+\mathbb{A}_{2,n})$
    where $\mathbb{A}_{1,n}$ and $\mathbb{A}_{2,n}$ are CSL algebras for each $n$.
  • (ii)
    If in addition $\mathcal{U}$ is self-adjoint, we can write it in the form
    (4) $\mathcal{U}=\bigcap_{n=1}^{\infty}(\mathbb{B}_{n}+\mathbb{B}_{n}^{*})$
    where $\mathbb{B}_{n}$ is a CSL algebra for each $n$.
Proof.
By Lemma 2.1, for ϕ:=assignitalic-$\phi:=$ Map$(\mathcal{U})$ and for a sequence of projections $P_{n}$ in $\mathbb{D}$ we can write
$\mathcal{U}=\bigcap_{n=1}^{\infty}\mathcal{U}_{n}$
where $\mathcal{U}_{n}=\{T\in B(\mathcal{H}):\ \phi(P_{n})^{\perp}TP_{n}=0\}$. Then, by Proposition 2.2 $(i)$, $\mathcal{U}_{n}=\mathbb{A}_{1,n}+\mathbb{A}_{2,n}$ for some CSL algebras $\mathbb{A}_{1,n}$, $\mathbb{A}_{2,n}$.
In case $\mathcal{U}$ is self-adjoint, we clearly have
$\mathcal{U}=\bigcap_{n=1}^{\infty}\Delta(\mathcal{U}_{n})$
and by Proposition 2.2 $(ii)$, for each $n$, we can write $\Delta(\mathcal{U}_{n})=\mathbb{B}_{n}+\mathbb{B}_{n}^{*}$ for some CSL algebra $\mathbb{B}_{n}$. ∎
Remark 2.1.
In case $\mathcal{H}$ is finite dimensional, the intersections (3) and (4) are clearly finite. Moreover even if $\mathcal{H}$ is not separable, by Lemma 2.1, we can still express $\mathcal{U}$ as
$\mathcal{U}=\bigcap_{P\in\mathbb{D}}\mathcal{U}_{P}$
where $\mathcal{U}_{P}=\{T\in B(\mathcal{H}):\ \phi(P)^{\perp}TP=0\}$, but this time the intersection might not be countable.
3. $C^{*}$-envelopes and the trivial intersection property
In this section, we introduce a property, called the trivial intersection property for concrete operator spaces and we show that the $C^{*}$-envelope of unital subspaces $\mathcal{U}\subseteq B(\mathcal{H})$ with this property is (*-isomorphic to) $C^{*}(\mathcal{U})$. In Theorem 3.3, we indicate some conditions, under which a space has the trivial intersection property and as we will see, Theorem 3.3 can be applied in order to compute the $C^{*}$-envelope in several different cases of unital operator spaces.
Definition 3.1.
We say that a closed subspace $\mathcal{U}\subseteq B(\mathcal{H})$ has the trivial intersection property if for every closed two-sided ideal $J$ of $C^{*}(\mathcal{U})$, $\mathcal{U}\cap J=\{0\}$ implies $J=\{0\}$.
In order to fix some notation, for two $C^{*}$-algebras $C,D$, by $C\cong D$ we mean that $C$ and $D$ are *-isomorphic. We denote by $\mathcal{K}(\mathcal{H})$ the set of all compact operators in $B(\mathcal{H})$ and for every subset $V\subseteq B(\mathcal{H})$ we set $\mathcal{K}(V):=\mathcal{K}(\mathcal{H})\cap V$.
Proposition 3.2.
If $\mathcal{U}\subseteq B(\mathcal{H})$ is a closed unital subspace that has the trivial intersection property, then $C^{*}_{e}(\mathcal{U})\cong C^{*}(\mathcal{U})$.
Proof.
Let $J$ be a boundary ideal of $\mathcal{U}$, i.e. the restriction $\pi:=q|_{\mathcal{U}}$ of the quotient map $q$ from $C^{*}(\mathcal{U})$ onto $C^{*}(\mathcal{U})/J$ is a complete isometry. In particular, $\pi$ is one-to-one, therefore $\mathcal{U}\cap J=\{0\}$, so by the trivial intersection property, $J=\{0\}$. This proves that the only boundary ideal of $\mathcal{U}$ is $\{0\}$, therefore $C^{*}_{e}(\mathcal{U})\cong C^{*}(\mathcal{U})$. ∎
Theorem 3.3.
Suppose that $\mathcal{U}\subseteq B(\mathcal{H})$ is a unital subspace and that there exists a subset $\mathcal{S}$ of $\mathcal{K}(C^{*}(\mathcal{U}))$ such that $\mathcal{S}C^{*}(\mathcal{U})^{\prime}\subseteq\mathcal{U}$ and $I\in\overline{[\mathcal{S}+\mathcal{S}^{*}]}^{WOT}$. Then, $\mathcal{U}$ has the trivial intersection property.
Proof.
Suppose $J$ is an ideal of $C^{*}(\mathcal{U})$ with $J\cap\mathcal{U}=\{0\}$ and $q$ is the quotient map from $C^{*}(\mathcal{U})$ onto $C^{*}(\mathcal{U})/J$. Then $\overline{J}^{w^{*}}$ is an ideal of $\mathbb{A}=\overline{C^{*}(\mathcal{U})}^{w^{*}}$ and the latter is a von Neumann algebra. By [Tak01, Proposition II.3.12], $\overline{J}^{w^{*}}=\mathbb{A}p$ for some projection $p$ in the center of $\mathbb{A}$. Clearly $p\in\overline{J}^{w^{*}}$, so by Kaplansky’s density Theorem, there exists a net $(a_{i})\subseteq J$ with $\|a_{i}\|\leq 1$ for each $i$, such that $a_{i}\to p$ strongly. It is then easy to see that $Ka_{i}\to Kp$ in the norm topology, for each $K\in\mathcal{S}$. Since the center of $\mathbb{A}$ is contained in $C^{*}(\mathcal{U})^{\prime}$, we have $Kp\in\mathcal{U}$. Therefore, using that $Ka_{i}\in J$, we see that $0=q(Ka_{i})\to q(Kp)$, thus $Kp=0$ for every $K\in\mathcal{S}$. Since $p\in C^{*}(\mathcal{U})^{\prime}$, we also have $\mathcal{S}^{*}p=0$, thus $p=0$ by our hypothesis that $I\in\overline{[\mathcal{S}+\mathcal{S}^{*}]}^{WOT}$ and $J=\{0\}$. ∎
Theorem 3.3 provides a large class of Corollaries/Examples of concrete unital operator spaces, with the trivial intersection property.
Corollary 3.4.
Suppose $\mathcal{U}\subseteq B(\mathcal{H})$ is a unital subspace such that $\overline{\mathcal{K}(\mathcal{U})+\mathcal{K}(\mathcal{U})^{*}}^{WOT}=B(\mathcal{H})$. Then $\mathcal{U}$ has the trivial intersection property.
Proof.
Since $\mathcal{K}(\mathcal{U})+\mathcal{K}(\mathcal{U})^{*}\subseteq C^{*}(\mathcal{U})$, we have $C^{*}(\mathcal{U})^{\prime}=\mathbb{C}I$. Thus, for $\mathcal{S}=\mathcal{K}(\mathcal{U})$, all conditions of Theorem 3.3 are met. ∎
A subspace $\mathcal{U}\subseteq B(\mathcal{H})$ is called strongly reflexive if there exists a set $\mathscr{R}$ of rank one operators such that $\mathcal{U}=$ $(\mathscr{R})$. We have the following:
Corollary 3.5.
Suppose $\mathcal{U}\subseteq B(\mathcal{H})$ is strongly reflexive and a unital $\mathbb{D}$-bimodule over a masa $\mathbb{D}\subseteq B(\mathcal{H})$. Then $\mathcal{U}$ has the trivial intersection property.
Proof.
It follows from [EKS98] that the subspace spanned by the rank one operators in $\mathcal{U}$ is WOT-dense in $\mathcal{U}$. Therefore, $\mathcal{U}=\overline{\mathcal{K}(\mathcal{U})}^{WOT}$. Thus, for $\mathcal{S}=\mathcal{K}(\mathcal{U})$, we have $\mathcal{S}C^{*}(\mathcal{U})^{\prime}\subseteq\mathcal{U}\mathbb{D}\subseteq\mathcal{U}$ and $I\in\overline{\mathcal{S}}^{WOT}$ hence, by Theorem 3.3, $\mathcal{U}$ has the trivial intersection property. ∎
Corollary 3.6.
Let $\mathcal{V}\subseteq B(\mathcal{H})$ be any set satisfying $\mathcal{V}^{\prime}\subseteq\mathcal{V}$ and $\mathcal{U}$ is a unital subspace with $\mathcal{U}\subseteq\overline{\mathcal{K}(\mathcal{U})+\mathcal{K}(\mathcal{U})^{*}}^{WOT}$ and $\mathcal{V}\mathcal{U}\subseteq\mathcal{U}$ ((((or $\mathcal{U}\mathcal{V}\subseteq\mathcal{U})$. Then $\mathcal{U}$ has the trivial intersection property.
Proof.
Clearly, $\mathcal{V}\subseteq\mathcal{U}$, therefore $C^{*}(\mathcal{U})^{\prime}\subseteq\mathcal{V}$. Then, if we set $\mathcal{S}=\mathcal{K}(\mathcal{U})$, we get $\mathcal{S}C^{*}(\mathcal{U})^{\prime}\subseteq\mathcal{U}$ (in either one of the cases $\mathcal{V}\mathcal{U}\subseteq\mathcal{U}$, $\mathcal{U}\mathcal{V}\subseteq\mathcal{U}$). Finally, $I\in\mathcal{U}\subseteq\overline{\mathcal{K}(\mathcal{U})+\mathcal{K}(\mathcal{U})^{*}}^{WOT}$, therefore we can apply Theorem 3.3. ∎
Example 3.7.
If $\mathcal{U}\subseteq B(\mathcal{H})$ is a unital $\mathbb{D}$-bimodule over some masa $\mathbb{D}\subseteq B(\mathcal{H})$, which satisfies $\mathcal{U}\subseteq\overline{\mathcal{K}(\mathcal{U})+\mathcal{K}(\mathcal{U})^{*}}^{WOT}$ then by Corollary 3.6, $\mathcal{U}$ has the trivial intersection property.
Example 3.8.
Let $\mathcal{L}$ be a CSL acting on a Hilbert space $\mathcal{H}$ and set $\mathcal{U}=\overline{\text{Alg}(\mathcal{L})+\mathcal{K}(\mathcal{H})}^{\|\cdot\|}$. Since $\mathcal{L}^{\prime}\subseteq$ $(\mathcal{L})$, we have $\mathcal{L}^{\prime}\mathcal{U}\subseteq\mathcal{U}$. Since $\mathcal{L}\subseteq\mathcal{L}^{\prime}$ and $\mathcal{K}(\mathcal{U})=\mathcal{K}(\mathcal{H})$, we can apply Corollary 3.6 with $\mathcal{V}=\mathcal{L}^{\prime}$.
For the last case, we refer the reader to [Dav88] for the definition of a completely distributive lattice.
Corollary 3.9.
Every completely distributive CSL algebra has the trivial intersection property.
Proof.
Let $\mathcal{U}=\textup{Alg}(\mathcal{L})$, where $\mathcal{L}$ is a completely distributive CSL. Then, by [Dav88, Theorem 23.7] we have $\mathcal{U}=\overline{\mathcal{K}(\mathcal{U})}^{w^{*}}$, so we can apply Corollary 3.6 with $\mathcal{V}=\mathcal{L}^{\prime}$. ∎
For the rest of the section, we deal with subspaces of $B(\ell^{2}(\Gamma))$, where $\Gamma$ is any set, which contain the discrete masa $\ell^{\infty}(\Gamma)$ and we shall prove the following:
Theorem 3.10.
Suppose $\mathcal{U}_{1},\mathcal{U}_{2}$ are unital operator subspaces of $B(\ell^{2}(\Gamma_{1}))$ and $B(\ell^{2}(\Gamma_{2}))$ respectively, and each $\mathcal{U}_{i}$ contains the masa $\mathbb{D}_{i}:=\ell^{\infty}(\Gamma_{i})$. Let $\theta:U_{1}\rightarrow U_{2}$ be a unital surjective complete isometry. Then, $\theta$ is a unitary equivalence, i.e. there exists a unitary $W:\ell^{2}(\Gamma_{1})\rightarrow\ell^{2}(\Gamma_{2})$ such that $\theta(T)=WTW^{*}$ for all $T\in U_{1}$.
For the proof, we need the following Lemma, which is a direct consequence of Theorem 3.3 and Proposition 3.2:
Lemma 3.11.
Suppose $\mathcal{U}$ is a unital operator space acting on $\mathcal{H}=\ell^{2}(\Gamma)$ that contains the masa $\mathbb{D}=\ell^{\infty}(\Gamma)$. Then, $C^{*}_{e}(\mathcal{U})\cong C^{*}(\mathcal{U})$.
Proof.
Let $\mathcal{S}=\mathcal{K}(\mathbb{D})\subseteq\mathcal{K}((C^{*}(\mathcal{U}))$. Since $C^{*}(\mathcal{U})^{\prime}\subseteq\mathbb{D}$, we have $C^{*}(\mathcal{U})^{\prime}\mathcal{S}\subseteq\mathbb{D}\subseteq\mathcal{U}$. Furthermore, $I\in\overline{\mathcal{S}}^{w^{*}}$, thus by Theorem 3.3, $\mathcal{U}$ has the trivial intersection property and then $C^{*}_{e}(\mathcal{U})\cong C^{*}(\mathcal{U})$ by Proposition 3.2. ∎
In what follows, for any subset $A\subseteq\Gamma$, we will write $P(A)$ for the projection corresponding to the element $(a_{\gamma})_{\gamma\in\Gamma}$, where $a_{i}=1$ if $i\in A$ and $a_{i}=0$ otherwise. Clearly $P(A)\in\ell^{\infty}(\Gamma)$ for every $A\subseteq\Gamma$.
Proof of Theorem 3.10.
We set $\mathbb{A}:=C^{*}(\mathcal{U}_{1})$ and $\mathbb{B}:=C^{*}(\mathcal{U}_{2})$. By Theorem 3.11, these are *-isomorphic to $C^{*}_{e}(\mathcal{U}_{1})$ and $C^{*}_{e}(\mathcal{U}_{2})$ respectively, hence there exists a *-isomorphism $\tilde{\theta}:\mathbb{A}\rightarrow\mathbb{B}$ that extends $\theta$. We observe that $\tilde{\theta}$ carries $\mathbb{A}^{\prime}$ onto $\mathbb{B}^{\prime}$. Since $\mathbb{A}^{\prime}\subseteq\ell^{\infty}(\Gamma_{1})$ and $\mathbb{B}^{\prime}\subseteq\ell^{\infty}(\Gamma_{2})$, it is well-known that there exists a unique family $\{Q_{i}:i\in\mathcal{I}\}$ of minimal orthogonal projections in $\mathbb{A}^{\prime}$ such that $\sum_{i}Q_{i}=I$. Since $\tilde{\theta}$ is an order isomorphism, it follows that $\{\tilde{\theta}(Q_{i}):i\in\mathcal{I}\}$ is a set of minimal orthogonal projections in $\mathbb{B}^{\prime}$ with $\sum_{i}\tilde{\theta}(Q_{i})=I$. It is then not so hard to see that
$\overline{\mathbb{A}}^{w^{*}}=\bigoplus_{i}Q_{i}B(\ell^{2}(\Gamma_{1}))Q_{i}$
$\overline{\mathbb{B}}^{w^{*}}=\bigoplus_{i}\tilde{\theta}(Q_{i})B(\ell^{2}(\Gamma_{2}))\tilde{\theta}(Q_{i}).$
For each $i$, we set $\mathbb{A}_{i}:=Q_{i}\mathbb{A}Q_{i}\subseteq\mathbb{A}$, $\mathbb{B}_{i}:=\tilde{\theta}(Q_{i})\mathbb{B}\tilde{\theta}(Q_{i})\subseteq\mathbb{B}$ and $\Theta(A_{i})$ for the unique subset of $\Gamma_{2}$ such that $\tilde{\theta}(Q_{i})=P(\Theta(A_{i}))$, so we can view $\mathbb{A}_{i}$ and $\mathbb{B}_{i}$ as $C^{*}$-subalgebras of $B(\ell^{2}(A_{i}))$ and $B(\ell^{2}(\Theta(A_{i}))$ respectively. We observe that
$\overline{\mathbb{A}_{i}}^{w^{*}}=Q_{i}\overline{\mathbb{A}}^{w^{*}}Q_{i}=B(\ell^{2}(A_{i}))$
and by the same reasoning $\overline{\mathbb{B}_{i}}^{w^{*}}=B(\ell^{2}(\Theta(A_{i})))$, so both $\mathbb{A}_{i}$ and $\mathbb{B}_{i}$ are irreducible. Moreover, since $\mathbb{A}_{i}$ and $\mathbb{B}_{i}$ contain $\ell^{\infty}(A_{i})$ and $\ell^{\infty}(\Theta(A_{i}))$ respectively, which in turn contain non-zero compact operators, by [Dav96, Corollary I.10.4], $\mathcal{K}(\ell^{2}(A_{i}))\subseteq\mathbb{A}_{i}$ and $\mathcal{K}(\ell^{2}(\Theta(A_{i})))\subseteq\mathbb{B}_{i}$. Thus, by [KR97, Theorem 10.4.6], there exists a unitary operator $W_{i}:\ell^{2}(A_{i})\rightarrow\ell^{2}(\Theta(A_{i}))$ such that
$\tilde{\theta}(T)=W_{i}TW_{i}^{*}$
for every $T\in\mathbb{A}_{i}$. Finally, if we define $W:=\oplus_{i}W_{i}:\ell^{2}(\Gamma_{1})\rightarrow\ell^{2}(\Gamma_{2})$, which is unitary, we get $\tilde{\theta}(T)=WTW^{*}$ for every $T\in\mathbb{A}$. ∎
Remark 3.1.
Suppose $\mathbb{A}\subseteq B(\ell^{2}(\Gamma_{1}))$, $\mathbb{B}\subseteq B(\ell^{2}(\Gamma_{2}))$ are *-isomorphic von Neumann algebras that contain the masas $\ell^{\infty}(\Gamma_{1})$ and $\ell^{\infty}(\Gamma_{2})$ respectively. Then, by Theorem 3.10, they are unitarily equivalent. In this case though, the unitary operator $U$ that induces the equivalence can be picked so that $\ell^{\infty}(\Gamma_{2})=U\ell^{\infty}(\Gamma_{1})U^{*}$. In order to see this, let $\theta:\mathbb{A}\rightarrow\mathbb{B}$ be a *-isomorphism. Then, following the proof of Theorem 3.10, we obtain an orthogonal family $\{Q_{i}:i\in\mathcal{I}\}$ of minimal projections in $\mathbb{A}^{\prime}$ such that $\sum_{i}Q_{i}=I$ and $\{\theta(Q_{i}):i\in\mathcal{I}\}$ is also an orthogonal family of minimal projections in $\mathbb{B}^{\prime}$ with $\sum_{i}\theta(Q_{i})=I$, and since $\mathbb{A}$, $\mathbb{B}$ are von Neumann algebras,
$\mathbb{A}=\bigoplus_{i}Q_{i}B(\ell^{2}(\Gamma_{1}))Q_{i}$
$\mathbb{B}=\bigoplus_{i}\theta(Q_{i})B(\ell^{2}(\Gamma_{2}))\theta(Q_{i}).$
For certain subsets $A_{i}\subseteq\Gamma_{1}$ and $B_{i}\subseteq\Gamma_{2}$, we can view $Q_{i}B(\ell^{2}(\Gamma_{1}))Q_{i}$ and $\theta(Q_{i})B(\ell^{2}(\Gamma_{2}))\theta(Q_{i})$ as $B(\ell^{2}(A_{i}))$ and $B(\ell^{2}(B_{i}))$ respectively, and again by the proof of Theorem 3.10, we get that $\ell^{2}(A_{i})$ and $\ell^{2}(B_{i})$ are isomorphic, therefore $A_{i}$ and $B_{i}$ have the same cardinality. Let $f_{i}:A_{i}\rightarrow B_{i}$ be one-to-one and onto. If $\{e_{\gamma}:\gamma\in\Gamma_{1}\}$ and $\{e_{\delta}:\delta\in\Gamma_{2}\}$ are the canonical bases of $\ell^{2}(\Gamma_{1})$ and $B(\ell^{2}(\Gamma_{2}))$ respectively, for each $i$ we set $U_{i}:\ell^{2}(A_{i})\rightarrow\ell^{2}(B_{i})$ by $U_{i}e_{\gamma}=e_{f_{i}(\gamma)}$, for each $\gamma\in A_{i}$. It is clear then that $\mathbb{B}=U\mathbb{A}U^{*}$ and $\ell^{\infty}(\Gamma_{2})=U\ell^{\infty}(\Gamma_{1})U^{*}$, where $U=\oplus_{i}U_{i}$, and that $Ue_{\gamma}=e_{f(\gamma)}$, where $f(\gamma)=f_{i}(\gamma)$ for the unique $i$ such that $\gamma\in A_{i}$.
4. Discrete Groups
In this section, we shall apply results of the previous section to discrete groups. For a group $G$, we view the elements of $\ell^{2}(G)$ as functions $f:G\rightarrow\mathbb{C}$. Therefore,
$\sum_{g\in G}|f(g)|^{2}<\infty$
for all $f\in\ell^{2}(G)$. The inner product in $\ell^{2}(G)$ is
$\langle f,h\rangle=\sum_{g\in G}f(g)\overline{h(g)}.$
We represent elements in $\ell^{\infty}(G)\subseteq B(\ell^{2}(G))$ in the form $M_{f}$, where $f:G\rightarrow\mathbb{C}$ is a bounded function and $M_{f}(g)=fg$ for all $g\in\ell^{2}(G)$. For two elements $f,h\in\ell^{2}(G)$, we write $f\otimes h^{*}$ for the rank one operator
$x\mapsto\langle x,h\rangle f.$
For all subsets $\kappa\subseteq G$, we write $\chi_{\kappa}$ for the characteristic function on $\kappa$, and $M_{\kappa}:=M_{\chi_{\kappa}}$ which is a projection in $\ell^{\infty}(G)$. For any subset $\Omega\subseteq G\times G$, we set
(5) $\mathbb{M}(\Omega):=\{T\in B(\ell^{2}(G)):M_{\lambda}TM_{\kappa}=0\ \text{for all}\ \kappa,\lambda\subseteq G\ \text{with}\ \kappa\times\lambda\subseteq\Omega^{c}\}.$
Clearly, $\mathbb{M}(\Omega)$ is a $w^{*}$-closed subspace of $B(\ell^{2}(G))$, which is also a $\ell^{\infty}(G)$-bimodule.
For every $g\in G$, we set $e_{g}:=\chi_{\{g\}}\in\ell^{2}(G)$. We observe the following:
Lemma 4.1.
For any $\Omega\subseteq G\times G$, we have
$\mathbb{M}(\Omega)=\overline{[e_{h}\otimes e_{g}^{*}:\ (g,h)\in\Omega]}^{w^{*}}.$
Proof.
Suppose $(g,h)\in\Omega$ and $\kappa\times\lambda\subseteq\Omega^{c}$. Let $x\in M_{\kappa}$, $y\in M_{\lambda}$. Then,
$\langle(e_{h}\otimes e_{g}^{*})x,y\rangle=\langle\langle x,e_{g}\rangle e_{h},y\rangle=\langle x,e_{g}\rangle\langle e_{h},y\rangle$
and at least one of these is zero, since $(g,h)\not\in\kappa\times\lambda$. Thus $M_{\lambda}(e_{h}\otimes e_{g}^{*})M_{\kappa}=0$ for all $\kappa\times\lambda\subseteq\Omega^{c}$, i.e. $e_{h}\otimes e_{g}^{*}\in\mathbb{M}(\Omega)$. Since $\mathbb{M}(\Omega)$ is a $w^{*}$-closed subspace, it follows that
$\overline{[e_{h}\otimes e_{g}^{*}:\ (g,h)\in\Omega]}^{w^{*}}\subseteq\mathbb{M}(\Omega).$
For the converse, let $T\in\mathbb{M}(\Omega)$. Since
$\sum_{g\in G}e_{g}\otimes e_{g}^{*}=I$
we have
$T=\sum_{g,h}(e_{h}\otimes e_{h}^{*})T(e_{g}\otimes e_{g}^{*})=\sum_{g,h}\langle Te_{g},e_{h}\rangle e_{h}\otimes e_{g}^{*}$
and since clearly $\langle Te_{g},e_{h}\rangle=0$ whenever $(g,h)\not\in\Omega$, the inverse inclusion holds as well.
Now, let $E\subseteq G$. We set $E^{\star}:=\{(g,h)\in G\times G:g^{-1}h\in E\}$. The following proposition relates $E$ with $\mathbb{M}(E^{\star})$ regarding their algebraic properties:
Proposition 4.2.
Suppose $G$ is a group with identity element $1$ and $E\subseteq G$. Then:
  • (i)
    $1\in E$ if and only if $\mathbb{M}(E^{\star})$ is unital.
  • (ii)
    $E=E^{-1}$ if and only if $\mathbb{M}(E^{\star})$ is self-adjoint.
  • (iii)
    $E\cdot E\subseteq E$ if and only if $\mathbb{M}(E^{\star})$ is an algebra.
In particular, $E$ is a subgroup of $G$ if and only if $\mathbb{M}(E^{\star})$ is a von Neumann algebra.
Proof.
  • (i)
    We have $1\in E$ if and only if $\{(g,g):g\in G\}\subseteq E^{\star}$. On the other hand, since $M_{\kappa}M_{\lambda}=0\Leftrightarrow\kappa\cap\lambda=\emptyset$, we have $I\in\mathbb{M}(E^{\star})$ if and only if $\kappa\cap\lambda=\emptyset$ for all $(\kappa\times\lambda)\cap E^{\star}=\emptyset$, which is equivalent to the fact $\{(g,g):g\in G\}\subseteq E^{\star}$.
  • (ii)
    $E=E^{-1}\Leftrightarrow$ $E^{\star}$ is symmetric, i.e. $(g,h)\in E^{\star}\Leftrightarrow(h,g)\in E^{\star}$. Thus $(\kappa\times\lambda)\cap E^{\star}=\emptyset$ if and only if $(\lambda\times\kappa)\cap E^{\star}=\emptyset$, which is equivalent to $\mathbb{M}(E^{\star})$ being self-adjoint.
  • (iii)
    Suppose $E\cdot E\subseteq E$. By Lemma 4.1, it suffices to show that if $(g_{1},h_{1})$, $(g_{2},h_{2})\in E^{\star}$, the operator $(e_{h_{1}}\otimes e_{g_{1}}^{*})(e_{h_{2}}\otimes e_{g_{2}}^{*})\in\mathbb{M}(E^{\star})$. Since
    $(e_{h_{1}}\otimes e_{g_{1}}^{*})(e_{h_{2}}\otimes e_{g_{2}}^{*})=\langle e_{h_{2}},e_{g_{1}}\rangle e_{h_{1}}\otimes e_{g_{2}}^{*}$
    it suffices to assume $g_{1}=h_{2}$. Since $g_{1}^{-1}h_{1}\in E$ and $g_{2}^{-1}h_{2}\in E$, we have
    $g_{2}^{-1}h_{2}g_{1}^{-1}h_{1}=g_{2}^{-1}h_{1}\in E$
    therefore $(g_{2},h_{1})\in E^{\star}$ and $e_{h_{1}}\otimes e_{g_{2}}^{*}\in\mathbb{M}(E^{\star})$. Now suppose that $\mathbb{M}(E^{\star})$ is an algebra. For $g,h\in E$, we have $(g^{-1},1)$, $(1,h)\in E^{\star}$, therefore $e_{1}\otimes e_{g^{-1}}^{*}$ and $e_{h}\otimes e_{1}^{*}$ are in $\mathbb{M}(E^{\star})$. Since the latter is an algebra,
    $e_{h}\otimes e_{g^{-1}}^{*}=(e_{h}\otimes e_{1}^{*})(e_{1}\otimes e_{g^{-1}}^{*})\in\mathbb{M}(E^{\star})$
    hence $(g^{-1},h)\in E^{\star}$, or $gh\in E$.
For every set $E\subseteq G$, we denote by $\langle E\rangle$ the subgroup generated by $E$, and for a subset $A\subseteq B(\ell^{2}(G))$, we denote by $W^{*}(A)$ the von Neumann algebra generated by $A$. The following is a corollary of Proposition 4.2:
Corollary 4.3.
For every nonempty set $E\subseteq G$, we have
$W^{*}(\mathbb{M}(E^{\star}))=\mathbb{M}(\langle E\rangle^{\star}).$
Proof.
By Proposition 4.2, $\mathbb{M}(\langle E\rangle^{\star})$ is a von Neumann algebra, which clearly contains $\mathbb{M}(E^{\star})$, therefore $W^{*}(\mathbb{M}(E^{\star}))\subseteq\mathbb{M}(\langle E\rangle^{\star})$. Now, by Lemma 4.1, it suffices to prove that for every $(g,h)\in\langle E\rangle^{\star}$, $e_{h}\otimes e_{g}^{*}\in W^{*}(\mathbb{M}(E^{\star}))$. We have $g^{-1}h\in\langle E\rangle$, so there exist $x_{1},..,x_{n}\in E\cup E^{-1}$ such that $g^{-1}h=x_{1}..x_{n}$ i.e. $h=gx_{1}..x_{n}$. We will show that $e_{gx_{1}..x_{k}}\otimes e_{g}^{*}\in W^{*}(\mathbb{M}(E^{\star}))$ for every $1\leq k\leq n$ by induction:
For $k=1$, we have $x_{1}\in E\cup E^{-1}$. In case $x_{1}\in E$, $(g,gx_{1})\in E^{\star}$, thus $e_{gx_{1}}\otimes e_{g}^{*}\in\mathbb{M}(E^{\star})$. In case $x_{1}\in E^{-1}$, $(gx_{1},g)\in E^{\star}$, which implies $e_{g}\otimes e_{gx_{1}}^{*}\in\mathbb{M}(E^{\star})$, thus $e_{gx_{1}}\otimes e_{g}^{*}\in W^{*}(\mathbb{M}(E^{\star}))$.
Now suppose we have $e_{gx_{1}..x_{k}}\otimes e_{g}^{*}\in W^{*}(\mathbb{M}(E^{\star}))$ for some $k<n$. We have $x_{k+1}\in E\cup E^{-1}$. In case $x_{k+1}\in E$, $(gx_{1}..x_{k},gx_{1}..x_{k}x_{k+1})\in E^{\star}$, therefore $e_{gx_{1}..x_{k}x_{k+1}}\otimes e_{gx_{1}..x_{k}}^{*}\in\mathbb{M}(E^{\star})$ and thus
$e_{gx_{1}..x_{k}x_{k+1}}\otimes e_{g}^{*}=(e_{gx_{1}..x_{k}x_{k+1}}\otimes e_{gx_{1}..x_{k}}^{*})(e_{gx_{1}..x_{k}}\otimes e_{g}^{*})\in W^{*}(\mathbb{M}(E^{\star})).$
If $x_{k+1}\in E^{-1}$, $(gx_{1}..x_{k}x_{k+1},gx_{1}..x_{k})\in E^{\star}$, so $e_{gx_{1}..x_{k}}\otimes e_{gx_{1}..x_{k}x_{k+1}}^{*}\in\mathbb{M}(E^{\star})$ and therefore
$e_{gx_{1}..x_{k}x_{k+1}}\otimes e_{g}^{*}=(e_{gx_{1}..x_{k}}\otimes e_{gx_{1}..x_{k}x_{k+1}}^{*})^{*}(e_{gx_{1}..x_{k}}\otimes e_{g}^{*})\in W^{*}(\mathbb{M}(E^{\star})).$
The proof is now complete. ∎
Theorem 4.4.
Suppose $H_{1}$, $H_{2}$ are subgroups of $G_{1}$ and $G_{2}$ respectively. The following are equivalent:
  • (i)
    $\mathbb{M}(H_{1}^{\star})$ and $\mathbb{M}(H_{2}^{\star})$ are unitarily equivalent.
  • (ii)
    $\mathbb{M}(H_{1}^{\star})$ and $\mathbb{M}(H_{2}^{\star})$ are *-isomorphic.
  • (iii)
    $H_{1}$ and $H_{2}$ have the same cardinality and the same indices in $G_{1}$, $G_{2}$ respectively, i.e. $|H_{1}|=|H_{2}|$ and $[G_{1}:H_{1}]=[G_{2}:H_{2}]$.
Proof.
Obviously, $(i)\Rightarrow(ii)$. Now suppose that $(ii)$ holds. By Remark 3.1, there exists a unitary operator $U:\ell^{2}(G_{1})\rightarrow\ell^{2}(G_{2})$ and a bijection $f:G_{1}\rightarrow G_{2}$ such that $\mathbb{M}(H_{2}^{\star})=U\mathbb{M}(H_{1}^{\star})U^{*}$ and $Ue_{g}=e_{f(g)}$ for each $g\in G_{1}$. We have the following:
$g^{-1}h\in H_{1}\iff(g,h)\in H_{1}^{\star}\iff e_{h}\otimes e_{g}^{*}\in\mathbb{M}(H_{1}^{\star})$
$\iff U(e_{h}\otimes e_{g}^{*})U^{*}\in\mathbb{M}(H_{2}^{\star})\iff e_{f(h)}\otimes e_{f(g)}^{*}\in\mathbb{M}(H_{2}^{\star})$
$\iff(f(g),f(h))\in H_{2}^{\star}\iff f(g)^{-1}f(h)\in H_{2}.$
Using this, we see that the map
$\tilde{f}:G_{1}/H_{1}\rightarrow G_{2}/H_{2}$
$gH_{1}\mapsto f(g)H_{2}$
is well-defined, one-to-one and onto. This means $[G_{1}:H_{1}]=[G_{2}:H_{2}]$. Also, by our earlier observation, $g\in H_{1}$ if and only if $f(1)^{-1}f(g)\in H_{2}$, that is $f(g)\in f(1)H_{2}$. Since $f$ is one-to-one and onto, we get $f(H_{1})=f(1)H_{2}$. Therefore,
$|H_{1}|=|f(H_{1})|=|f(1)H_{2}|=|H_{2}|$
and we have proved $(ii)\Rightarrow(iii)$.
Finally, suppose that $|H_{1}|=|H_{2}|$ and $[G_{1}:H_{1}]=[G_{2}:H_{2}]$. Then, there exists a bijective map $f:G_{1}\rightarrow G_{2}$ such that $f(gH_{1})=f(g)H_{2}$ for every $g\in G_{1}$. This implies $g^{-1}h\in H_{1}\iff f(g)^{-1}f(h)\in H_{2}$ for all $g,h\in G_{1}$. Now, if we set $U:\ell^{2}(G_{1})\rightarrow\ell^{2}(G_{2})$ by $Ue_{g}=e_{f(g)}$, using Lemma 4.1 one easily observes that $\mathbb{M}(H_{2}^{\star})=U\mathbb{M}(H_{1}^{\star})U^{*}$, therefore $(iii)\Rightarrow(i)$. ∎
A direct Corollary using Lagrange’s Theorem is the following:
Corollary 4.5.
Suppose $G$ is a finite group and $H_{1}$, $H_{2}$ are subgroups of $G$. The following are equivalent:
  • (i)
    $\mathbb{M}(H_{1}^{\star})$ and $\mathbb{M}(H_{2}^{\star})$ are unitarily equivalent.
  • (ii)
    $\mathbb{M}(H_{1}^{\star})$ and $\mathbb{M}(H_{2}^{\star})$ are *-isomorphic.
  • (iii)
    $|H_{1}|=|H_{2}|$.
Remark 4.1.
In case $G$ is finite and $H_{1}$, $H_{2}$ are two subgroups isomorphic to each other, by Corollary 4.5, we clearly have $\mathbb{M}(H_{1}^{\star})\cong\mathbb{M}(H_{2}^{\star})$. However, this is not true in general for infinite groups. For example, if $G=\mathbb{Z}$, $H_{1}=\mathbb{Z}$, $H_{2}=2\mathbb{Z}$, the groups $H_{1}$, $H_{2}$ are isomorphic, but $[G:H_{1}]=1$, $[G:H_{2}]=2$, so by Theorem 4.4, $\mathbb{M}(H_{1}^{\star})\not\cong\mathbb{M}(H_{2}^{\star})$. Also, $\mathbb{M}(H_{1}^{\star})\cong\mathbb{M}(H_{2}^{\star})$ does not mean in general that $H_{1}$ and $H_{2}$ are isomorphic (even in the finite case). For example, for $G=\mathbb{Z}_{2}\times\mathbb{Z}_{4}$, the subgroups $H_{1}=\langle(0,1)\rangle$, $H_{2}=\langle(0,2),(1,0)\rangle$ are not isomorphic, but they both have 4 elements, therefore by Corollary 4.5, $\mathbb{M}(E_{1}^{\star})\cong\mathbb{M}(E_{2}^{\star})$.
Corollary 4.6.
Let $E_{1}$, $E_{2}$ be subsets of the groups $G_{1}$, $G_{2}$ respectively, both containing the identities, and suppose $\mathbb{M}(E_{1}^{\star})$ and $\mathbb{M}(E_{2}^{\star})$ are completely isometrically isomorphic via a unital map. Then, $|\langle E_{1}\rangle|=|\langle E_{2}\rangle|$ and $[G_{1}:\langle E_{1}\rangle]=[G_{2}:\langle E_{2}\rangle]$.
Proof.
By Proposition 4.2, each $\mathbb{M}(E_{i}^{\star})$ contains the identity, and since it is a $\ell^{\infty}(G_{i})$-bimodule, it contains $\ell^{\infty}(G_{i})$. Therefore, by (the proof of) Theorem 3.10, the $C^{*}$-algebras $C^{*}(\mathbb{M}(E_{1}^{\star}))$, $C^{*}(\mathbb{M}(E_{2}^{\star}))$ are unitarily equivalent, hence so are their $w^{*}$-closures, i.e. $W^{*}(\mathbb{M}(E_{1}^{\star}))$ and $W^{*}(\mathbb{M}(E_{2}^{\star}))$. By Corollary 4.3, we get $\mathbb{M}(\langle E_{1}\rangle^{\star})\cong\mathbb{M}(\langle E_{2}\rangle^{*})$ and the conclusion follows from Theorem 4.4. ∎
Acknowledgment
The author would like to thank Professor G. K. Eleftherakis for his great guidance and his suggestions for this note, as well as Professor A. Katavolos for his very useful observations.
Declarations
Data availability statement Data sharing is not applicable to this article as no data sets were generated or analyzed during this study.
Conflict of interest The author declares that there are no conflicts of interest.